OpenStudy (amistre64):

spose i wanted to calculate the probability of getting 2 of a kind in a 5 card poker hand ....

OpenStudy (amistre64):

would it be a good idea to do this: first card can be any card: 1/13 next card assume the pair: 3/51 next 3 cards different: 48/50, 47/49, 46/48

OpenStudy (amistre64):

then multiply that by the number of ways to get 2 of a kind in a hand of 5?

OpenStudy (amistre64):

ive gotta run to class, so ill read this later if yall want to argue amongst yourselves :)

OpenStudy (anonymous):

I looked this up on the internet, maybe you should read it,

OpenStudy (anonymous):

The probability of getting two SPECIFIC jacks is the probablity of getting one of them, multiplied by the conditional probability of getting the other one given that the first one was also received, multiplied by the conditional probability of getting no more jacks other than the first two.. The first jack can go to either one of the 5 cards in the hand or to one of the other 47 spots, The probability is 5/52. If the first jack is known to be in the hand, there are only 4 spots left in the hand and only 51 cards unaccounted for, so the conditional probablity of getting the second jack is 4/51. With 50 cards remaining, and two of them jacks, the proability of getting three non-jacks out of the three remaining cards in the hand is (48/50)(47/49)(46/48). Finally, there are 6 distinct combinations of two out of the four jacks, (Think of the 4 sides of a square plus the two diagonals, or just calculate C(4,2).) So, there are 6 different mutually-exclusive ways, each with the above probability product, to get a hand with a pair of jacks. These combined to: p = 6(5/52)(4/51)(48/50)(47/49)(46/48) which is about 4%. If all you wanted was the probability of getting a hand with exactly two jacks (a pair) then that's your answer. If you want a hand that is called "a pair of jacks", then you have to rule out a two-pair hand or a full house. (3 or 4 of a kind in jacks have already been eliminated, as have been straights and flushes.) To do this, modify the product above to: p = 6(5/52)(4/51)(48/50)(44/49)(40/48) because after picking the first non jack (48 choices out of the 50 remaining cards), you cant pick that card in another suit, so the number of choices for the 2nd non-jack character has been reduced to 44 cards (52 cards - 4 jacks - 4 cards of the 3rd card rank) out of the 49 cards unaccounted for. The final product is now 40/48 for the same reason since the final card can't be jack (-4 cards), can't match the third card (-4 more cards) for (52-4-4-4=40) desirable cards out of the 48 remaining. That's about 3 and a quarter percent

OpenStudy (dan815):

@amistre64 you are not accounting for the 3 or 4 of a kind case

OpenStudy (amistre64):

at the moment, i am just considering getting a hand that has ONLY 2 of a kind; since 2 of a kind is valued differently as a hand then 3 of a kind, or 4 of a kind ... even a full house has 2 of a kind built into it.

OpenStudy (amistre64):

i had consider a possible binomial setup like this; which gives me about 4%, which is what my original idea turned up as well \[\large \binom 52~(\frac1{13})^2~(\frac{12}{13})^3\]

OpenStudy (dan815):

i dunno maybe im doing it wrong lol

OpenStudy (dan815):

the probability of getting atleast a 2 pair is way too high for my stuff

OpenStudy (dan815):

like 60-70%

OpenStudy (dan815):

lemme think about another method then

OpenStudy (dan815):

4-5% for 2 pair is about right though

OpenStudy (amistre64):

"at least" 2 of a kind is around 50% if i trust the internet :)

OpenStudy (dan815):